Write down the equations for the real electric and magnetic fields for a monochromatic plane wave of amplitude E(subscript 0), frequency omega, and phase angle phi=0,that is travelling in the direction from the origin to the point (1,1,1), with polarisation parallel to the xz plane.
Answer #1

The general equation of electric field of a Plane wave is given by

E(r,t)=Eo(r,t)(ei(K.r-ωt+φ)) (here i signifies √-1 )

r denotes position vector

B(r,t) = (K x E)/ω ( x denotes vector cross product)

where K is the wave vector of propagation.

vector along the propagation direction form (0,0,0) to (1,1,1)
K =(k/√3 )(x^+y^+z^ )

Polarised along the KZ plane means the electric field oscillates in the KZ plane

∴ E(r,t)= (Eoxx^+Eozz^)ei(K.r-ωt)

(Eox)2+(Eoy)2= Eo2

K.r =( k/√3 )(x+y+z)

Real electric field is given by

E = (Eoxx^+Eozz^)Cos(( k/√3 )(x+y+z)-ωt)

Real magnetic field is given by

B =(k/ω√3) (Eozx^+(E0x -Eoz)y^ + Eoxz^)Cos(( k/√3)(x+y+z)-ωt)

answered by: Shone